Convert to weights

How much would a chrome book weigh in quarters?

How much would it weigh in $1 bills?

How much would it weigh in $100 bills?

How many of it would it be needed to pay off the debt of $22 trillion?
1
Expert's answer
2020-06-21T16:41:30-0400

If the chromebook weights for example 1.1 kg and a quarter - 5.67 g, 1$ bill=1 g then a chromebook will weight

  • 194 in quarters (quarter - 5.67 g)
  • 1100 in 1$ bills (1$ bill=1 g)
  • 1375 in 100$ (100$=0.8 g)
  • 160 million (if the debt will be paid in 100$ bills, the mass of $22 trillion will be: 0.8*22*10^(12)/100=176*109 g;  amount of it needed to pay off the debt: 176*109/1100= 160 million chromebooks);